### v2

# Automatically detecting scientific political science texts from a large general document index

Abstract
This technical report outlines the filtering approach applied to the collection of the Bielefeld Academic Search Engine (BASE) data to extract articles from the political science domain. We combined hard and soft filters to address entries with different available metadata, e.g. title, abstract or keywords. The hard filter is a weighted keyword-based approach. The soft filter uses a multilingual BERT-based classification model, trained to detect scientific articles from the political science domain. We evaluated both approaches using an annotated dataset, consisting of scientific articles from different scientific domains. The weighted keyword-based approach achieved the highest total accuracy of 0.88. The multilingual BERT-based classification model was fine-tuned using a dataset of 14,178 abstracts from scientific articles and reached the highest total accuracy of 0.98. The proposed filtering approach can be applied for filtering metadata from other scientific domains and therefore improve the overview of the domain-related literature and facilitate efficiency in research.
Keywords Classification  $\cdot$ Political science  $\cdot$ Pollux  $\cdot$ Python
1 Introduction
The specialized information service (FID) Political Science - Pollux  [^1] https://www.pollux-fid.de/ aims to offer political scientists subject-specific and fast access to special literature and information relevant to their research. Pollux obtains data from different providers, one of them is BASE  [^2] https://www.base-search.net/ (Bielefeld Academic Search Engine), which indexes the metadata of all kinds of academically relevant resources. Many BASE entries have no domain classification, therefore we developed a filtering approach which can detect entries pertinent to political science from the huge collection of BASE metadata. Additionally, publications relevant to political science also might be included in other scientific domains. State and University Library Bremen (SuUB) has used BASE filtering methods for their Electronic library (E-LIB) since 2012 (Blenkle et al., 2016; Blenkle and Nölte, 2017). The core BASE filtering approach including a weighted keyword-based approach was developed in 2016 by the SuUB team for the Pollux project.
In this technical report, we describe the training procedure, training data and usage of two classification models which are part of one of the modules (soft filter) of our filtering approach (see Figure 2). Additionally, we provide an evaluation of both classification models and a weighted keyword-based approach (hard filter) on a specially designed test dataset.
1.1 BASE
The BASE is a search engine for academic web resources operated by Bielefeld University Library. The BASE provides over 340 million documents from different content providers and consists of various kinds of academic resources, e.g., journals, books digital collections, etc  [^3] https://www.base-search.net/about/en/index.php?. Many BASE entries are classified by the Dewey Decimal Classification (DDC). The DDC is a system for organizing library contents by dividing all knowledge into ten categories, each assigned a range of 100 numbers. Aside from DDC BASE data has many other important metadata, such as source, journal name, authors’ names, journal name, etc. Figure 1 shows the BASE base interface and basic metadata.
Figure 1: Example of the BASE interface and metadata
1.2 Filtering workflow
Developing approaches for filtering political science articles from a collection of unlabeled or multidisciplinary scientific articles is crucial for several reasons. Hołyst et al. (2024) argue that excessive information constrains our ability to assess information and make optimal decisions. Therefore, excluding articles irrelevant to political science can enhance political scientists’ time efficiency and enable focusing on the materials most pertinent to their work. Political science often intersects with other disciplines. Finding articles related to politics in other disciplines will provide political researchers with opportunities for interdisciplinary collaboration, enrich their perspectives and enhance the depth of their research. Political science is a dynamic field, with new events, policies, and trends continuously emerging. Timely access to relevant literature allows political researchers to remain abreast of the latest developments in the political field, enabling informed analyses and discussions.
The filtering approach for the BASE data collection comprises different modules to address entries with different metadata available as illustrated in Figure 2. Initially, the data is filtered according to parameters such as data type and source. Subsequently, in the second filter articles already categorized under the DDC system as political science (DDC 320-328) are selected. The third filter selects articles with a valid abstract. For articles lacking abstracts, a keyword-based filter (hard filter) is applied. Articles containing abstracts undergo a language filter, and those written in permissible languages are then subjected to a BERT-based classification model (soft filter).
Figure 2: Pipeline for the BASE filtering approach
The two major modules of this approach are so-called hard and soft filters. The hard filter is a weighted keyword-based filter approach. We developed a set of keywords which we believe to be effective for identifying data related to the political science domain. Each keyword was manually assigned with a score. The score is assigned according to the keyword’s relevancy for the domain, as Table 1 demonstrates. Figure 3 shows a simplified workflow for the soft filter. The filtering algorithm looks for the keywords in the metadata and returns a sum of scores of all found keywords. If this score is more or equal to 1 the article is marked as relevant to the domain. This module is applied to the entities which do not have a full abstract, and therefore the filter is applied to the entity’s title and keywords if the latter are available.
Figure 3: Soft filter simplified workflow
keyword score
*politik 1
bürgerkrieg 0.6
policy 0.4
Table 1: Example of a keyword list for the hard filter
The soft filter uses a BERT-based classification model, trained to detect scientific articles from the political science domain. The model was finetuned on abstracts from scientific articles, therefore the filter is applied to the entity’s title and abstract.
2 Soft Filter
For our soft-filter approach, we trained two classification models: English and multilingual, which can detect texts from the political science domain.
2.1 English classification model
The English classification model to detect texts from the political science domain (SSciBERT_politics) [^4] https://huggingface.co/kalawinka/SSciBERT_politics is based on the SSCI-SciBERT (Shen et al., 2023) model. SSCI-SciBERT was trained on the abstracts of social science research articles, therefore, we opted to use this model. The model was fine-tuned using a dataset of 2,919 abstracts from scientific articles retrieved from the BASE and POLLUX collections.
The SSciBERT_politics was initially integrated into the soft filter of the BASE filtering pipeline in the POLLUX infrastructure. Nevertheless, the model can be used separately with the Transformers library  [^5] https://huggingface.co/docs/transformers/en/index, as Listing 1 demonstrates.
1from transformers import AutoTokenizer, AutoModelForSequenceClassification
2from transformers import pipeline
3
4tokenizer = AutoTokenizer.from_pretrained(’kalawinka/SSciBERT_politics’)
5model = AutoModelForSequenceClassification.from_pretrained(’kalawinka/SSciBERT_politics’)
6pipe = pipeline("text-classification", model=model, tokenizer = tokenizer, max_length=512, truncation=True)
7
8pipe(’add scientific abstract’)
Listing 1: Usage example of SSciBERT_politics with the Transformers library
Applying this model to the abstract of the technical report yields the output presented in Listing 2.
1 [{’label’: ’multi’, ’score’: 0.9991981387138367}]
Listing 2: Example of SSciBERT_politics output
The model was fine-tuned utilizing the Transformers library and using the AdamW optimizer [^6] https://huggingface.co/docs/bitsandbytes/main/en/reference/optim/adamw. Evaluation was done at the end of each epoch. The total number of training epochs to perform was set to 4. The initial learning rate was set to 5e-5. The training was performed using one NVIDIA A40 GPU.
Listing 3 demonstrates the training parameters used to fine-tune the SSciBERT_politics model.
1training_args = TrainingArguments(output_dir=output_dir,
2 evaluation_strategy="epoch",
3 num_train_epochs = 4,
4 )
Listing 3: Training parameters for the finetuning of SSciBERT_politics with the Transformers library
2.1.1 Training data
Due to a lack of manually annotated data, we employed a semi-automated method to generate the labelled training dataset. Figure 4-A illustrates the distribution of articles across the training, test, and validation corpora. Accordingly, the test and validation datasets each contain 20% of the data, while the training corpus comprises 60% of the total data. Data from the BASE and POLLUX collections were utilized to construct the training corpus. To acquire scientific abstracts within the domain of political science, we utilized both POLLUX and BASE collections. We compiled a list of political science journals in English. Afterwards, we collected articles from the POLLUX collection published in these journals containing an abstract and published after 2018. Following, we collected articles from the BASE collection adhering to specific criteria. Only articles with the types article or review, DDC classification 320-328 (Political science), written in English, containing an abstract and published after 2018 was selected. Figure 4-C shows the distribution of sources in the ’politics’ category. To collect data from other (not political science) domains, we utilized only articles from the BASE collection. We sampled articles which correspond to the following criteria. Only articles with the types article or review, DDC classification not 320-328 (Political science), written in English containing an abstract and published after 2015 were selected. Additionally, we designed our query to restrict the selection only to articles which do not contain the keyword ’politi*’ in title, keywords or abstracts. Figure 4-B shows the distribution of scientific domains in the ’multi’ category [^7] One article can belong to multiple scientific domains.. After collecting the data from both collections, we verified the language of the abstracts to ensure they were in English and excluded abstracts containing fewer than 20 words.
Figure 4: Distribution of training data in English corpus
2.2 The multilingual classification model
The multilingual classification model to detect texts from the political science domain (bert-base-ml-politics) [^8] https://huggingface.co/kalawinka/bert-base-ml-politics is based on BERT multilingual base model (Devlin et al., 2018). The BERT multilingual base model was trained on multilingual Wikipedia data (102 languages). As SSCI-SciBERT was trained only for English it was not applicable for the fine-tuning with multilingual data. Therefore, we utilized a multilingual language representation model for the fine-tuning. The model was fine-tuned using a dataset of 14,178 abstracts from scientific articles retrieved from the BASE and POLLUX collections of scientific articles. Abstracts from scientific articles in 3 languages (English, German and French) were selected for the training.
The model is integrated into the soft filter of the BASE filtering pipeline in the POLLUX infrastructure. Nevertheless, the model can be used separately with the Transformers library, as Listing 4 demonstrates.
1from transformers import AutoTokenizer, AutoModelForSequenceClassification
2from transformers import pipeline
3
4tokenizer = AutoTokenizer.from_pretrained(’kalawinka/bert-base-ml-politics’)
5model = AutoModelForSequenceClassification.from_pretrained(’kalawinka/bert-base-ml-politics’)
6pipe = pipeline("text-classification", model=model, tokenizer = tokenizer, max_length=512, truncation=True)
7
8pipe(’add scientific abstract’)
Listing 4: Usage example of bert-base-ml-politics with the Transformers library
Applying this model to the abstract of this technical report produces the output illustrated in Listing 5.
1 [{’label’: ’politics’, ’score’: 0.9972042441368103}]
Listing 5: Example of bert-base-ml-politics output
As SSciBERT_politics, the model was finetuned using the Transformers library as Listing 6 demonstrates. The AdamW algorithm was applied to minimise the loss function. Evaluation was done at the end of each epoch. The total number of training epochs to perform was set to 3. The initial learning rate was set to 5e-5. The training was performed using one NVIDIA A40 GPU. Training time comprises 52,20334 minutes.
1training_args = TrainingArguments(output_dir=output_dir,
2 evaluation_strategy="epoch",
3 num_train_epochs = 3,
4 )
Listing 6: Training parameters for the finetuning of bert-base-ml-politics with the Transformers library
2.2.1 Training data
The collection of the training data was similar to the one described in Section 2.1.1. Data from the BASE and POLLUX collections were utilized to construct the training corpus. A semi-automated method was employed to generate the labelled training dataset. To obtain scientific abstracts from the political science domain we used both POLLUX and BASE collection. We created a list of political science journals from German-, English- and French-speaking countries. Afterwards, we collected articles from the POLLUX collection published in these journals containing an abstract and publication date. In the next step, we collected articles from the BASE collection corresponding to the following criteria. Only articles with the types article or review, DDC classification 320-328 (Political science), written in English, German or French, containing an abstract were selected. Figure 5-D shows the distribution of sources in the ’politics’ category. To collect data from other (not political science) domains, we utilized only articles from the BASE collection. We sampled articles which correspond to the following criteria. Only articles with the types article or review, DDC classification not 320-328 (Political science), written in English, German or French, containing an abstract were selected Additionally we designed our query to restrict the selection only to articles which do not contain the keyword ’politi*’ in title, keywords or abstracts. Figure 5 -C shows the distribution of scientific domains in the ’multi’ category [^9] One article can belong to multiple scientific domains.. After collecting the data from both collections, we verified the language of the abstracts to ensure they were in English, German, or French and excluded abstracts containing fewer than 20 words.
Figure 5-A shows the distribution of articles in the training, test and validation corpora. Thus, test and validation datasets contain 20% each and the training corpus comprises 60% of all data. Figure 5-B shows the distribution of languages in the training data. English is a prevailing language, as we encountered difficulties locating sufficient data in German and French. Furthermore, our training data is limited to these languages because we were unable to find adequate training data in other languages that met our criteria. Specifically, we required data that included both a title and a valid abstract. Many multilingual entries contained abstracts only in English. Additionally, for the BASE dataset, we required entries with the DDC classification.
Figure 5: Distribution of training data in the multilingual corpus
3 Evaluation
We evaluated both approaches using an annotated dataset of 4,726 abstracts from different scientific domains. The weighted keyword-based approach was tested in two modes: applied to title, keywords and abstract, and applied only to title and keywords. The SSciBERT_politics model was evaluated on a separate dataset consisting of 973 abstracts written in English.
Table 2 shows an evaluation of the approaches by class for all languages. Keyword-based approach, applied to title, abstract and keywords showed a high accuracy of 0,87, as well as a high F1-score for both classes: multi and politics. The keyword-based approach applied only to titles and keywords showed a slightly deteriorating performance of 0,8. Class politics has high precision and low recall, at the same time class multi has high recall and low precision. Both BERT-based classifiers showed high total accuracy and F1-scores for all classes. The bert-base-ml-politics showed higher accuracy (0.978) than the SSciBERT_politics model (0.9).
approach label precision recall f1-score support accuracy
keyword filter (title, abstract, keywords) politics 0.871 0.846 0.859 2143
keyword filter (title, abstract, keywords) multi 0.876 0.896 0.886 2583 0.874
keyword filter (title, keywords) politics 0.947 0.593 0.729 2143
keyword filter (title, keywords) multi 0.742 0.973 0.842 2583 0.800
bert-base-ml-politics politics 0.975 0.978 0.976 2143
bert-base-ml-politics multi 0.981 0.979 0.980 2583 0.978
SSciBERT_politics politics 0.889 0.902 0.895 460
SSciBERT_politics multi 0.911 0.899 0.905 513 0.900
Table 2: Evaluation by class
Table 3 presents the evaluation results by class and language, which align with the findings in Table Table 2. The bert-base-ml-politics model demonstrates the best performance across all languages. Conversely, the keyword-based approach applied to titles and keywords exhibits low performance for French, with an F1-score of 0.343. However, French had the least amount of training and evaluation data within the entire corpus, potentially introducing bias into the evaluation for all algorithms
approach label language precision recall f1-score support
keyword filter (title, abstract, keywords) politics English 0.931 0.839 0.883 1212
keyword filter (title, abstract, keywords) multi English 0.848 0.936 0.890 1164
keyword filter (title, abstract, keywords) politics German 0.843 0.861 0.852 783
keyword filter (title, abstract, keywords) multi German 0.856 0.838 0.847 776
keyword filter (title, abstract, keywords) politics French 0.647 0.831 0.728 148
keyword filter (title, abstract, keywords) multi French 0.958 0.896 0.926 643
keyword filter (title, keywords) politics English 0.976 0.568 0.718 1212
keyword filter (title, keywords) multi English 0.686 0.985 0.809 1164
keyword filter (title, keywords) politics German 0.943 0.699 0.803 783
keyword filter (title, keywords) multi German 0.759 0.957 0.847 776
keyword filter (title, keywords) politics French 0.625 0.236 0.343 148
keyword filter (title, keywords) multi French 0.846 0.967 0.903 643
bert-base-ml-politics politics English 0.989 0.993 0.991 1212
bert-base-ml-politics multi English 0.992 0.989 0.991 1164
bert-base-ml-politics politics German 0.952 0.958 0.955 783
bert-base-ml-politics multi German 0.957 0.951 0.954 776
bert-base-ml-politics politics French 0.979 0.959 0.969 148
bert-base-ml-politics multi French 0.991 0.995 0.993 643
Table 3: Evaluation by class and language
4 Conclusion
In general, the BERT-based classifier demonstrated the highest accuracy, but at the same time the highest processing time, approximately 4.5 seconds (on a CPU) to process a single abstract. Conversely, the keyword filter, although slightly less accurate, exhibited significantly faster performance compared to the classification model (less than 1 second to process a single abstract). The efficacy of the keyword filter improves when applied to all available metadata.
Both methodologies are applicable to filtering political science data. The keyword-based approach is particularly advantageous when dealing with large datasets and limited computational resources for running complex models. Furthermore, this approach is also suitable for processing articles with incomplete metadata.
The main limitation of the method is a lack of manually annotated training data. To address this issue we applied a semi-automated labelling approach. The proposed filtering approach can be applied for filtering metadata from other scientific domains and therefore improve the overview of the domain-related literature and facilitate efficiency in research.
Acknowledgements
Nina Smirnova acknowledges support by Deutsche Forschungsgemeinschaft (DFG) under grant number MA 3964/7-2, the Fachinformationsdienst Politikwissenschaft – Pollux. The core BASE-filtering pipeline was developed by the team of the State and University Library Bremen (SuUB): Martin Blenkle, Elmar Haake, Thore Christiansen, Marie-Saphira Flug, Lena Klaproth. The weighted keyword-based filtering approach was developed by Daniel Opitz. Michael Czolkoß-Hettwer contributed to the creation of the list of keywords for the political science domain. Nina Smirnova was founded by the European Union under the Horizon Europe OMINO - Overcoming Multilevel Information Overload (101086321, https://ominoproject.eu/). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them.